OpenStudy (anonymous):

if the roots of the equation ax^2 + bx + c =0 are of the form x/(x-1) and (x+1)/x then the value of (a+b+c)^2 is?

OpenStudy (anonymous):

\[(a + b+c)^2 = a^2 + b^2 + c^2 + 2(ab + bc + ca)\] Will this help or not?

OpenStudy (shubhamsrg):

@experimentX ,,product of roots is not 1..maybe your mis-read..

OpenStudy (experimentx):

Oh .. i misread.

OpenStudy (shubhamsrg):

product of roots = (x+1)/(x-1) = c/a sum of roots = (2x^2 - 1)(x^2 -1) = -b/a => (c+b)/a = (x+1)/(x-1) - (2x^2 - 1)(x^2 -1) = something.. i dont know how to eliminate a now..

OpenStudy (shubhamsrg):

as c+b = f(a)

OpenStudy (anonymous):

Sum Of Roots: -b/a And product of Roots = c/a

OpenStudy (anonymous):

\[r1*r2=\frac{x+1}{x-1}\] \[\frac{c}{a}=\frac{x+1}{x-1}\] use componendo and dividendo to find x.....

OpenStudy (experimentx):

seems we can add it too ..

OpenStudy (anonymous):

\[\frac{x}{x-1} + \frac{x+1}{x} = \frac{-b}{a}\]

OpenStudy (anonymous):

\[= \frac{x^2 + x^2 - 1}{x(x-1)} = \frac{2x^2 - 1}{x(x-1)} = \frac{-b}{a}\]

OpenStudy (experimentx):

we still need one equation.

OpenStudy (shubhamsrg):

i got it..i min i'll post it

OpenStudy (shubhamsrg):

i mean 1 min**

OpenStudy (shubhamsrg):

product of roots = (x+1)/(x-1) = c/a => x= (c+a)/(c-a) on substitutions, we see roots are : (c+a)/2a and 2c/(a+c) now sum of roots = -b/a = [(a+c)^2 + 4ac ]/(2a(a+c) = -b/a => a^2 + c^2 +6ac = -2ab -2bc =>(a+b+c)^2 = b^2 - 4ac = discriminant .. oops,,maybe you want it in terms of x only ?

OpenStudy (shubhamsrg):

can you tell me the ans if you have it @Yahoo!

OpenStudy (anonymous):

b^2-4ac

OpenStudy (shubhamsrg):

i mean the book's ans ?

OpenStudy (anonymous):

yup man u r correct!!!

OpenStudy (shubhamsrg):

hmm,,phew!! :P